OpenStudy (anonymous):

(x+6)/(x^2-4x-21) + (x+2)/(x^2-6x-27)

OpenStudy (anonymous):

I have (x+6)/(x-7)(x+4) + (x+2)/(x-9)(x+3) but not sure what to do next

OpenStudy (anonymous):

(x+6)/(x-7)(x+4) sure of this one ?:

OpenStudy (anonymous):

i meant (x-7)(x+3)

OpenStudy (anonymous):

now, all you can do is to add them ! after "putting" them at the same denominator :)

OpenStudy (anonymous):

what about the numerator

OpenStudy (anonymous):

by multiplying the first "the numerator and the denominator by (x-9) and the second "the numerator and the denominator by (x-7) do this and see what you'll get !

OpenStudy (anonymous):

all i got was codes when you wrote that

OpenStudy (anonymous):

\[\frac{(x+6)(x-9)}{(x-7)(x+3)(x-9)}+\frac{(x+2)(x-7)}{(x-7)(x+3)(x-9)}\] now can you add them ?!

OpenStudy (anonymous):

what do you mean add them?

OpenStudy (anonymous):

like foil?

OpenStudy (anonymous):

\[\frac{a}{d}+\frac{b}{d}=???\]

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

:'( ! you don't know that \[\frac{a}{d}+\frac{b}{d}=\frac{a+b}{d}\] ??

OpenStudy (anonymous):

i understand but i dont know how to do it with (x+6)(x-9) + (x-2)(x-7)

OpenStudy (anonymous):

so ! we agree that what you typed is equal to \[\frac{(x+6)(x-9)+(x+2)(x-7)}{(x-7)(x+3)(x-9)}\] ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so ! can you calculate (x+6)(x-9)+(x+2)(x-7) ?!

OpenStudy (anonymous):

no i already said that i dont know how to

OpenStudy (anonymous):

(x+6)(x-9)=(x)(x)+(x)(-9)+(6)(x)+(6)(-9)=x^2-3x-54 can you do the same for (x+2)(x-7) ?

OpenStudy (anonymous):

yeah i asked if you just foil and yous aid no

OpenStudy (anonymous):

but i get it now...thank you!

OpenStudy (anonymous):

I didn't say no :) !

OpenStudy (anonymous):

you're welcome ! anytime !